# LoRAverse: A Submodular Framework to Retrieve Diverse Adapters for Diffusion Models

### v1

@@ Abstract
**footnotetext: This work was conducted while the author was a research intern at Virginia Tech.
@@ 1 Introduction
LoRA [35] models have emerged as a powerful technique for model personalization. They can efficiently fine-tune pre-trained diffusion models without the need for extensive retraining or significant computational resources. They are designed to optimize low-rank, factorized weight matrices specifically for the attention layers and are typically used in conjunction with personalization methods such as DreamBooth [34] to generate personalized content on specific objects, individuals, or artistic styles. Since their introduction, LoRA models have gained widespread popularity among researchers, developers, and artists [11, 14]. Platforms like Civit.ai [2], which host over 100,000 LoRA models, have become major hubs for creators looking to extend the capabilities of pre-trained models to generate highly specific image styles, themes, or artistic directions. This flexibility allows users to target precise modifications without retraining the entire model, enabling fast adaptation and diverse creative outputs.
@@
Despite the vast potential, the sheer volume of available LoRA adapters makes it challenging for users to navigate and effectively use these models. Users often struggle to explore, evaluate, and combine different LoRA adapters to achieve desired outcomes. Without clear guidance or an efficient system for selecting the right combination, users are left with a trial-and-error approach that is both time-consuming and inefficient. Consider a user who wants to generate fantasy world concept designs with themes like ”ancient ruins,” ”futuristic cities,” and ”enchanted forests” using LoRA adapters. Users familiar with platforms like Civit.ai face an overwhelming choice among 100,000+ models, each varying in style, quality, and effectiveness. The challenge lies in selecting the best-fit adapter for their concept amid the massive range and diversity of LoRAs. Existing retrieval methods [26] use top-$K$ ranking and LLM selection. While relevant, top-$K$ ranking often surfaces similar models, leading to redundant recommendations and overlapping features, and relying on LLMs to prune these further limits the variation and biases the input prompt.
@@
In this paper, we explore the challenge of selecting the most relevant and diverse LoRA adapters from the vast database. We approach the task of finding suitable LoRA adapters, each tuned to generate highly specific styles and concepts, as a combinatorial optimization problem. Our objective is to identify a subset of LoRA adapters that are not only relevant but also diverse in the types of concepts they represent. This is crucial as many adapters may offer overlapping functionalities. Our objective function incorporates diversity by penalizing the selection of additional models from already-represented LoRA concepts. In order to achieve this, we cluster similar LoRA models, allowing our algorithm to prioritize unexplored clusters, thus enhancing the diversity of the selected subset. Our framework, LoRAverse, then ensures that selecting a LoRA from a cluster that has not yet been explored yields a higher gain. We formulate this task as a monotone submodular function maximization, for which there is a simple greedy algorithm guarantees that the solution obtained is $\epsilon$-close to the best possible solution [20].
@@ 2 Related Works
LoRA models provide an efficient solution for fine-tuning models on specific tasks with minimal parameter updates, achieving performance comparable to full fine-tuning while reducing computational and storage demands [10, 15, 17]. Their adaptability and efficiency have led to widespread adoption within open-source communities [41]. In the image domain, these models excel at combining diverse attributes–such as styles, characters, poses, and actions–to generate high-quality images that align with user prompts [23, 44]. However, with the abundance of available adapters, selecting the most relevant ones for specific generative tasks remains a significant challenge [25]. Retrieval-based methods like retrieval-augmented generation (RAG) address similar challenges by embedding text inputs into vectors and ranking results based on semantic similarity to the user prompt [8, 12, 19, 21, 31, 32, 26]. While effective, these methods often struggle to introduce diversity in the retrieved outputs. Submodular optimization, widely used for selecting diverse subsets from larger pools, provides an effective solution to this problem [20]. In particular, submodular techniques model the trade-off between relevance and diversity in adapter selection, ensuring that both highly relevant adapters and diverse image styles are maintained. Clustering methods, such as those used in GAN-based models to group semantically similar elements [6, 30, 37], further complement this by grouping similar adapters, enabling the selection of diverse yet relevant adapters for high-quality image generation. Recent diffusion‑based sampling strategies also tackle diversity, but at generation time rather than retrieval [7].
@@ 3 Background
Let $\mathcal{V}$ denote a set of elements $\mathcal{V}=\{v_{1},\ldots v_{n}\}$ (the ground set), and let $\mathcal{F}:2^{\mathcal{V}}\to\mathbb{R}$ represent a function that assigns a real value to any subset $\mathcal{P}\subseteq\mathcal{V}$. The goal is then to select a small subset $\mathcal{P}^{*}\subseteq\mathcal{V}$ of size $|\mathcal{P}|\leq n$ that maximizes the function $\mathcal{F}(\mathcal{P})$, i.e., $\mathcal{P}^{*}\in\operatorname*{arg\,max}_{\mathcal{P}\subseteq\mathcal{V},\,|\mathcal{P}|\leq n}\mathcal{F}(\mathcal{P})$. Solving this problem is intractable in general, but it has been shown that a greedy algorithm can be used to solve this equation almost optimally with an approximation factor of (11/e1-1/e), under the condition that the function $\mathcal{F}$ is monotone, non-decreasing, and submodular [38]. Starting with an empty set $\mathcal{P}_{0}=\emptyset$, the algorithm iteratively selects the element $v_{i}$ that maximizes the marginal gain: $v_{i}\in\operatorname*{arg\,max}_{v\in\mathcal{V}\setminus\mathcal{P}_{i-1}}\mathcal{F}(\mathcal{P}_{i-1}\cup\{v\})-\mathcal{F}(\mathcal{P}_{i-1})$. After $n$ iterations, the resulting subset $\mathcal{P}^{*}$ satisfies the following inequality: $\mathcal{F}(\mathcal{P}^{*})\geq(1-1/e)\mathcal{F}(\mathcal{P_{\text{opt}}})\approx 0.63\,\mathcal{F}(\mathcal{P}_{\text{opt}})$ where $\mathcal{P}_{\text{opt}}$ is the optimal subset. More formally, submodularity is defined as follows:
@@
The function $\mathcal{F}$ is called submodular if for all subsets $R\subseteq P\subseteq\mathcal{V}$ and any elements $v\in\mathcal{V}\setminus\mathcal{P}$, the following inequality holds: $\mathcal{F}(\mathcal{P}\cup\{v\})-\mathcal{F}(\mathcal{P})\leq\mathcal{F}(\mathcal{R}\cup\{v\})-\mathcal{F}(\mathcal{R})$. This form of submodularity directly satisfies the diminishing returns property: the value of adding $v$ never becomes larger as the context becomes larger [29].
@@ 4.2 Submodular Retriever
In our work, we treat user prompts as combinations of distinct concepts (§ 4.1) and aim to retrieve the top-$K$ most relevant and diverse LoRA models based on these concepts. A simple approach would be retrieving adapters using cosine similarity between the full prompt embedding and adapter embeddings. While this ensures relevance, it often results in redundancy, as retrieved adapters show only minor variations, limiting diversity (see Fig. 8). On the other hand, selecting adapters based solely on individual concept embeddings can introduce adapters relevant to specific concepts but misaligned with the overall prompt meaning (see Fig. 9). To address these issues, our submodular retriever module maximizes a monotone submodular function, balancing relevance and diversity. Using a greedy algorithm, we find an almost-optimal solution [20], ensuring the selected adapters are both relevant and diverse. Formally, let $\mathcal{A}$ be the set of available LoRA models, and we select a diverse, representative subset $\mathcal{P}\subseteq\mathcal{A}$, where relevance is based on the similarity between the prompt embedding and LoRA model embeddings from a Vision Language Model (VLM). We define the relevance objective as:
@@
However, this relevance-only function neglects diversity, often favoring multiple models of the same style (e.g., photorealistic) and overlooking others [2]. To address this, a common approach applies a diversity regularization to the objective function [20, 37], rewarding items selected from different clusters such that:
@@
The clustering objective ensures selecting models from unexplored clusters yields higher gains than from already-covered ones, balancing prompt relevance and diversity. To avoid the inefficiency of clustering the entire LoRA set, we first retrieve a subset relevant to the prompt using cosine similarity, then apply HDBSCAN [27] on their textual embeddings via BERTopic [13] (see SM § 15 for details). This groups LoRA models semantically, ensuring diverse and relevant selection. Then, the overall objective combines both:
@@
Finding the optimal subset is computationally infeasible, but a greedy algorithm can approximate the solution within a factor of 11/e1-1/e for monotone, non-negative, submodular functions [29]. In practice, these algorithms often surpass this theoretical bound, proving both efficient and reliable [20]. The submodular retriever brings this approach together by selecting relevant and diverse LoRA models for each concept while aligning with the full prompt. It clusters similar models and applies submodular function maximization. Formally, the submodular retriever $\mathcal{R}$ takes a set of concepts $\mathcal{T}(s)$ from the concept extractor and retrieves the adapter set $\mathcal{P}^{*}$:
@@ 5 Experiments
We compare LoRAverse with two baseline models: Stylus [26] and Stable-Diffusion-v1.5 (SD v1.5) [33], using the two of the most prominent model checkpoints Realistic-Vision-v6 and Counterfeit-v3. We used StylusDocsv2 [26] database, which includes around 75,000 LoRA models from platforms like CivitAI [2] and Hugging Face [1]. Images are generated via the Stable Diffusion Web UI [4] with 35 denoising steps and the DPM Solver++ scheduler [24]. During the experiments, we observed occasional style drift caused by certain LoRA models, which we mitigated by appending a debias prompt [^1] The debias prompt is ”realistic, high quality” for Realistic-Vision-v6 and ”anime style, high quality” for Counterfeit-v3 to the user input, ensuring style consistency—similar to Stylus. Additionally, to maintain the ethical integrity of LoRAverse, we implemented an adapter safety checker using gpt-4o (see SM § 13 for the full prompt) to filter out adapters containing inappropriate or explicit content. This ensures that the retrieved adapters align with professional expectations and ethical standards. In our submodular algorithm, we set the relevance and diversity trade-off parameters, $\lambda_{1}$ and $\lambda_{2}$, to 7.0 and 1.0, respectively, balancing both factors in the selection process.
@@ 5.1 Qualitative Experiments
We compare LoRAverse qualitatively with the baseline approaches in Fig. 3. Each comparison includes 15 images generated by our method, Stylus [26], and SD v1.5 [33], showing that our approach consistently produces the most diverse image sets. In contrast, the baseline methods often struggle to retrieve adapters that produce images with a wide range of diversity in aspects such as style, background, or primary content. While Stylus provides greater diversity than SD v1.5, its reliance on adapter retrieval based solely on cosine similarity leads to redundancy among the selected adapters (§ 5.3). Furthermore, its use of an LLM called composer [26] for selection introduces an inherent bias toward the input prompt, limiting the diversity of its outputs. Consequently, it cannot surpass a certain threshold in generating varied images. Our framework overcomes these limitations with a submodular-based adapter selection process, driven entirely by an algorithmic framework. For example, given the prompt ”A fantastical witch in a forest in vibrant style,” other methods struggle to generate witches with distinct clothing or diverse settings. In contrast, our method effectively captures and enhances this diversity, delivering a richer and more varied visual representation. Additional qualitative results and further examples are provided in SM § 9.
@@ 5.2 Quantitative Experiments
Image Diversity. We evaluate LoRAverse quantitatively using four benchmarks: CLIP (ViT-L/14) [16], Truncated CLIP Entropy (TCE) [18], Truncated Inception Entropy (TIE) [18], and pairwise image similarity (I2I). CLIP measures text-image alignment, while TCE and TIE assess diversity–TCE captures semantic diversity in CLIP’s text-image embedding space, and TIE quantifies visual diversity using InceptionV3 [39]. Additionally, I2I measures pairwise similarity using image-to-image CLIP score.
Our evaluation includes 15,000 images–5,000 per method and 20 per prompt–from MS-COCO [22] and PartiPrompts [43] using the Realistic-Vision-v6 checkpoint. We retrieve top-200 LoRA models per concept and selected top-8 of them using our method. As shown in Table 1, LoRAverse significantly enhances diversity while maintaining strong text-image alignment. Despite a 3.1% decrease in CLIP score, it compensates with a 16.5% increase in TCE and a 5.1% increase in TIE, demonstrating superior diversity. Additionally, its outputs are 7.3% more dissimilar than those of baseline methods, as reflected in I2I scores. Further quantitative experiments can be found in SM § 10 and 11.
LoRAverse preserves diversity across varying Classifier-Free Guidance (CFG) values (Fig. 4) and shifts the Pareto curve toward greater semantic and visual diversity, demonstrating its effectiveness. This confirms that our submodular-based approach outperforms LLM-based methods, like Stylus, in promoting diversity, making LoRAverse more versatile. The slight drop in CLIP score is expected, as increased diversity naturally broadens the output distribution, leading to greater variation in image composition. CLIP favors strict prompt adherence over generative diversity due to its training on large-scale web image-text pairs, which reinforce common associations over novel representations [3, 36]. Thus, the lower CLIP score reflects LoRAverse’s improved diversity rather than a decline in text-image alignment.
@@
VLM-as-a-Judge. We adopt the VLM-as-a-Judge approach to evaluate the image quality, textual alignment, and diversity of LoRAverse, Stylus, and SD v1.5, generating five images per prompt across 300 images with positions randomly shuffled to minimize bias [44]. We ask gpt-4o to rate images from 0 (poor) to 2 (high) for quality, alignment, and diversity. As shown in Fig. 5, LoRAverse demonstrates a clear advantage in diversity, significantly outperforming both Stylus and SD v1.5. It also maintains strong performance in textual alignment and visual quality, closely matching or surpassing the baselines. These results highlight LoRAverse’s ability to balance diversity and quality. Full prompt details and settings are provided in SM § 14.

### v2

@@ Abstract
[…]
@@ 1 Introduction
LoRA [ryu2023low] models have emerged as a powerful technique for model personalization. They can efficiently fine-tune pre-trained diffusion models without the need for extensive retraining or significant computational resources. They are designed to optimize low-rank, factorized weight matrices specifically for the attention layers and are typically used in conjunction with personalization methods such as DreamBooth [ruiz2023dreambooth] to generate personalized content on specific objects, individuals, or artistic styles. Since their introduction, LoRA models have gained widespread popularity among researchers, developers, and artists [gandikota2023concept, guo2023animatediff]. Platforms like Civit.ai [civitai_website], which host over 100,000 LoRA models, have become major hubs for creators looking to extend the capabilities of pre-trained models to generate highly specific image styles, themes, or artistic directions. This flexibility allows users to target precise modifications without retraining the entire model, enabling fast adaptation and diverse creative outputs.
@@
Despite the vast potential, the sheer volume of available LoRA adapters makes it challenging for users to navigate and effectively use these models. Users often struggle to explore, evaluate, and combine different LoRA adapters to achieve desired outcomes. Without clear guidance or an efficient system for selecting the right combination, users are left with a trial-and-error approach that is both time-consuming and inefficient. Consider a user who wants to generate fantasy world concept designs with themes like ”ancient ruins,” ”futuristic cities,” and ”enchanted forests” using LoRA adapters. Users familiar with platforms like Civit.ai face an overwhelming choice among 100,000+ models, each varying in style, quality, and effectiveness. The challenge lies in selecting the best-fit adapter for their concept amid the massive range and diversity of LoRAs. Existing retrieval methods [luo2024stylus] use top-$K$ ranking and LLM selection. While relevant, top-$K$ ranking often surfaces similar models, leading to redundant recommendations and overlapping features, and relying on LLMs to prune these further limits the variation and biases the input prompt.
@@
In this paper, we explore the challenge of selecting the most relevant and diverse LoRA adapters from the vast database. We approach the task of finding suitable LoRA adapters, each tuned to generate highly specific styles and concepts, as a combinatorial optimization problem. Our objective is to identify a subset of LoRA adapters that are not only relevant but also diverse in the types of concepts they represent. This is crucial as many adapters may offer overlapping functionalities. Our objective function incorporates diversity by penalizing the selection of additional models from already-represented LoRA concepts. In order to achieve this, we cluster similar LoRA models, allowing our algorithm to prioritize unexplored clusters, thus enhancing the diversity of the selected subset. Our framework, LoRAverse, then ensures that selecting a LoRA from a cluster that has not yet been explored yields a higher gain. We formulate this task as a monotone submodular function maximization, for which there is a simple greedy algorithm guarantees that the solution obtained is $\epsilon$-close to the best possible solution [lin2011class].
@@ 2 Related Works
LoRA models provide an efficient solution for fine-tuning models on specific tasks with minimal parameter updates, achieving performance comparable to full fine-tuning while reducing computational and storage demands [gal2022imageworthwordpersonalizing, ha2016hypernetworks, hu2021loralowrankadaptationlarge]. Their adaptability and efficiency have led to widespread adoption within open-source communities [wolf2020huggingfacestransformersstateoftheartnatural]. In the image domain, these models excel at combining diverse attributes–such as styles, characters, poses, and actions–to generate high-quality images that align with user prompts [liu2023unsupervisedcompositionalconceptsdiscovery, zhong2024multiloracompositionimagegeneration]. However, with the abundance of available adapters, selecting the most relevant ones for specific generative tasks remains a significant challenge [Modelverse2023]. Retrieval-based methods like retrieval-augmented generation (RAG) address similar challenges by embedding text inputs into vectors and ranking results based on semantic similarity to the user prompt [devlin2019bertpretrainingdeepbidirectional, gao2024retrievalaugmentedgenerationlargelanguage, lewis2020retrieval, lin2023vectorsearchopenaiembeddings, radford2021learning, reimers2019sentencebertsentenceembeddingsusing, luo2024stylus]. While effective, these methods often struggle to introduce diversity in the retrieved outputs. Submodular optimization, widely used for selecting diverse subsets from larger pools, provides an effective solution to this problem [lin2011class]. In particular, submodular techniques model the trade-off between relevance and diversity in adapter selection, ensuring that both highly relevant adapters and diverse image styles are maintained. Clustering methods, such as those used in GAN-based models to group semantically similar elements [collins2020editingstyleuncoveringlocal, pakhomov2021segmentationstyleunsupervisedsemantic, simsar2022fantasticstylechannelsthem], further complement this by grouping similar adapters, enabling the selection of diverse yet relevant adapters for high-quality image generation. Recent diffusion‑based sampling strategies also tackle diversity, but at generation time rather than retrieval [corso2023particleguidancenoniiddiverse].
@@ 3 Background
Let $\mathcal{V}$ denote a set of elements $\mathcal{V}=\{v_{1},\ldots v_{n}\}$ (the ground set), and let $\mathcal{F}:2^{\mathcal{V}}\to\mathbb{R}$ represent a function that assigns a real value to any subset $\mathcal{P}\subseteq\mathcal{V}$. The goal is then to select a small subset $\mathcal{P}^{*}\subseteq\mathcal{V}$ of size $|\mathcal{P}|\leq n$ that maximizes the function $\mathcal{F}(\mathcal{P})$, i.e., $\mathcal{P}^{*}\in\operatorname*{arg\,max}_{\mathcal{P}\subseteq\mathcal{V},\,|\mathcal{P}|\leq n}\mathcal{F}(\mathcal{P})$. Solving this problem is intractable in general, but it has been shown that a greedy algorithm can be used to solve this equation almost optimally with an approximation factor of (11/e1-1/e), under the condition that the function $\mathcal{F}$ is monotone, non-decreasing, and submodular [sviridenko2004note]. Starting with an empty set $\mathcal{P}_{0}=\emptyset$, the algorithm iteratively selects the element $v_{i}$ that maximizes the marginal gain: $v_{i}\in\operatorname*{arg\,max}_{v\in\mathcal{V}\setminus\mathcal{P}_{i-1}}\mathcal{F}(\mathcal{P}_{i-1}\cup\{v\})-\mathcal{F}(\mathcal{P}_{i-1})$. After $n$ iterations, the resulting subset $\mathcal{P}^{*}$ satisfies the following inequality: $\mathcal{F}(\mathcal{P}^{*})\geq(1-1/e)\mathcal{F}(\mathcal{P_{\text{opt}}})\approx 0.63\,\mathcal{F}(\mathcal{P}_{\text{opt}})$ where $\mathcal{P}_{\text{opt}}$ is the optimal subset. More formally, submodularity is defined as follows:
@@
The function $\mathcal{F}$ is called submodular if for all subsets $R\subseteq P\subseteq\mathcal{V}$ and any elements $v\in\mathcal{V}\setminus\mathcal{P}$, the following inequality holds: $\mathcal{F}(\mathcal{P}\cup\{v\})-\mathcal{F}(\mathcal{P})\leq\mathcal{F}(\mathcal{R}\cup\{v\})-\mathcal{F}(\mathcal{R})$. This form of submodularity directly satisfies the diminishing returns property: the value of adding $v$ never becomes larger as the context becomes larger [nemhauser1978analysis].
@@ 4.2 Submodular Retriever
In our work, we treat user prompts as combinations of distinct concepts (§ 4.1) and aim to retrieve the top-$K$ most relevant and diverse LoRA models based on these concepts. A simple approach would be retrieving adapters using cosine similarity between the full prompt embedding and adapter embeddings. While this ensures relevance, it often results in redundancy, as retrieved adapters show only minor variations, limiting diversity (see Fig. 8). On the other hand, selecting adapters based solely on individual concept embeddings can introduce adapters relevant to specific concepts but misaligned with the overall prompt meaning (see Fig. 9). To address these issues, our submodular retriever module maximizes a monotone submodular function, balancing relevance and diversity. Using a greedy algorithm, we find an almost-optimal solution [lin2011class], ensuring the selected adapters are both relevant and diverse. Formally, let $\mathcal{A}$ be the set of available LoRA models, and we select a diverse, representative subset $\mathcal{P}\subseteq\mathcal{A}$, where relevance is based on the similarity between the prompt embedding and LoRA model embeddings from a Vision Language Model (VLM). We define the relevance objective as:
@@
However, this relevance-only function neglects diversity, often favoring multiple models of the same style (e.g., photorealistic) and overlooking others [civitai_website]. To address this, a common approach applies a diversity regularization to the objective function [lin2011class, simsar2022fantasticstylechannelsthem], rewarding items selected from different clusters such that:
@@
The clustering objective ensures selecting models from unexplored clusters yields higher gains than from already-covered ones, balancing prompt relevance and diversity. To avoid the inefficiency of clustering the entire LoRA set, we first retrieve a subset relevant to the prompt using cosine similarity, then apply HDBSCAN [Malzer_2020] on their textual embeddings via BERTopic [grootendorst2022bertopic] (see SM § 15 for details). This groups LoRA models semantically, ensuring diverse and relevant selection. Then, the overall objective combines both:
@@
Finding the optimal subset is computationally infeasible, but a greedy algorithm can approximate the solution within a factor of 11/e1-1/e for monotone, non-negative, submodular functions [nemhauser1978analysis]. In practice, these algorithms often surpass this theoretical bound, proving both efficient and reliable [lin2011class]. The submodular retriever brings this approach together by selecting relevant and diverse LoRA models for each concept while aligning with the full prompt. It clusters similar models and applies submodular function maximization. Formally, the submodular retriever $\mathcal{R}$ takes a set of concepts $\mathcal{T}(s)$ from the concept extractor and retrieves the adapter set $\mathcal{P}^{*}$:
@@ 5 Experiments
We compare LoRAverse with two baseline models: Stylus [luo2024stylus] and Stable-Diffusion-v1.5 (SD v1.5) [rombach2022high], using the two of the most prominent model checkpoints Realistic-Vision-v6 and Counterfeit-v3. We used StylusDocsv2 [luo2024stylus] database, which includes around 75,000 LoRA models from platforms like CivitAI [civitai_website] and Hugging Face [hf_website]. Images are generated via the Stable Diffusion Web UI [AUTOMATIC1111_Stable_Diffusion_Web_2022] with 35 denoising steps and the DPM Solver++ scheduler [lu2023dpmsolverfastsolverguided]. During the experiments, we observed occasional style drift caused by certain LoRA models, which we mitigated by appending a debias prompt [^1] The debias prompt is ”realistic, high quality” for Realistic-Vision-v6 and ”anime style, high quality” for Counterfeit-v3 to the user input, ensuring style consistency—similar to Stylus. Additionally, to maintain the ethical integrity of LoRAverse, we implemented an adapter safety checker using gpt-4o (see SM § 13 for the full prompt) to filter out adapters containing inappropriate or explicit content. This ensures that the retrieved adapters align with professional expectations and ethical standards. In our submodular algorithm, we set the relevance and diversity trade-off parameters, $\lambda_{1}$ and $\lambda_{2}$, to 7.0 and 1.0, respectively, balancing both factors in the selection process.
@@ 5.1 Qualitative Experiments
We compare LoRAverse qualitatively with the baseline approaches in Fig. 3. Each comparison includes 15 images generated by our method, Stylus [luo2024stylus], and SD v1.5 [rombach2022high], showing that our approach consistently produces the most diverse image sets. In contrast, the baseline methods often struggle to retrieve adapters that produce images with a wide range of diversity in aspects such as style, background, or primary content. While Stylus provides greater diversity than SD v1.5, its reliance on adapter retrieval based solely on cosine similarity leads to redundancy among the selected adapters (§ 5.3). Furthermore, its use of an LLM called composer [luo2024stylus] for selection introduces an inherent bias toward the input prompt, limiting the diversity of its outputs. Consequently, it cannot surpass a certain threshold in generating varied images. Our framework overcomes these limitations with a submodular-based adapter selection process, driven entirely by an algorithmic framework. For example, given the prompt ”A fantastical witch in a forest in vibrant style,” other methods struggle to generate witches with distinct clothing or diverse settings. In contrast, our method effectively captures and enhances this diversity, delivering a richer and more varied visual representation. Additional qualitative results and further examples are provided in SM § 9.
@@ 5.2 Quantitative Experiments
Image Diversity. We evaluate LoRAverse quantitatively using four benchmarks: CLIP (ViT-L/14) [CLIPScore], Truncated CLIP Entropy (TCE) [ibarrola2024measuringdiversitycocreativeimage], Truncated Inception Entropy (TIE) [ibarrola2024measuringdiversitycocreativeimage], and pairwise image similarity (I2I). CLIP measures text-image alignment, while TCE and TIE assess diversity–TCE captures semantic diversity in CLIP’s text-image embedding space, and TIE quantifies visual diversity using InceptionV3 [szegedy2015rethinkinginceptionarchitecturecomputer]. Additionally, I2I measures pairwise similarity using image-to-image CLIP score.
Our evaluation includes 15,000 images–5,000 per method and 20 per prompt–from MS-COCO [lin2015microsoftcococommonobjects] and PartiPrompts [yu2022parti] using the Realistic-Vision-v6 checkpoint. We retrieve top-200 LoRA models per concept and selected top-8 of them using our method. As shown in Table 1, LoRAverse significantly enhances diversity while maintaining strong text-image alignment. Despite a 3.1% decrease in CLIP score, it compensates with a 16.5% increase in TCE and a 5.1% increase in TIE, demonstrating superior diversity. Additionally, its outputs are 7.3% more dissimilar than those of baseline methods, as reflected in I2I scores. Further quantitative experiments can be found in SM § 10 and 11.
LoRAverse preserves diversity across varying Classifier-Free Guidance (CFG) values (Fig. 4) and shifts the Pareto curve toward greater semantic and visual diversity, demonstrating its effectiveness. This confirms that our submodular-based approach outperforms LLM-based methods, like Stylus, in promoting diversity, making LoRAverse more versatile. The slight drop in CLIP score is expected, as increased diversity naturally broadens the output distribution, leading to greater variation in image composition. CLIP favors strict prompt adherence over generative diversity due to its training on large-scale web image-text pairs, which reinforce common associations over novel representations [alabdulmohsin2024clipbiasusefulbalancing, shao2023investigatinglimitationclipmodels]. Thus, the lower CLIP score reflects LoRAverse’s improved diversity rather than a decline in text-image alignment.
@@
VLM-as-a-Judge. We adopt the VLM-as-a-Judge approach to evaluate the image quality, textual alignment, and diversity of LoRAverse, Stylus, and SD v1.5, generating five images per prompt across 300 images with positions randomly shuffled to minimize bias [zhong2024multiloracompositionimagegeneration]. We ask gpt-4o to rate images from 0 (poor) to 2 (high) for quality, alignment, and diversity. As shown in Fig. 5, LoRAverse demonstrates a clear advantage in diversity, significantly outperforming both Stylus and SD v1.5. It also maintains strong performance in textual alignment and visual quality, closely matching or surpassing the baselines. These results highlight LoRAverse’s ability to balance diversity and quality. Full prompt details and settings are provided in SM § 14.